OpenStudy (anonymous):

I am having troubIe with college algebra. The question is (x+52)^(2/5)=144

OpenStudy (anonymous):

\[((x+52)^{2/5})^{5/2}=(144)^{5/2}\] \[(x+52)=(144)^{5/2}\]

OpenStudy (anonymous):

x+52=12^5